Status Report

Mission Status Report: GeneSat-1 Status Report: 19 Dec 2006 12 noon PST (Mission Day 4; Biology Experiment Day 2)

By SpaceRef Editor
December 20, 2006
Filed under , ,
Mission Status Report: GeneSat-1 Status Report: 19 Dec 2006 12 noon PST (Mission Day 4; Biology Experiment Day 2)

We continue to get Excellent Performance and Data from the GeneSat-1 Spacecraft. Status updates below from Mission Manager (Bruce Yost) and Chief Technologist (Tony  Ricco). Summary Status and ata examples can be seen at:

John Hines, GeneSat-1 Project Manager

John, and all,

The results to this point are nothing short of SPECTACULAR!! All the subsystems appear to be performing flawlessly. Optics are making stable measurements, fluidics have fed the E. coli, the bugs are growing well and GLOWING, temperature, pressure, and humidity are stable at the right values.

The E. coli began showing growth, from the light scattering measurements, about 2.8 hr after the valve opened to introduce the nutrients. By about 10 hr after feeding, 9 of the 10 bio-wells showed measurable growth. (A range of growth rates is expected because there are 2 different E. coli strains, and the exact number of bacteria that initiate the growth varies from well to well).

By 9 hours after feeding, the fastest growing well showed the 1st expression of green fluorescent protein. By about 20 hr after growth start, 6 of 10 biowells showed definite green fluorescence, 2 more appeared to be just edging above baseline. [To be confirmed by Matthew or Macarena: the remaining 2 biowells contain non-fluorescent controls, so little or no fluoresc is to be expected from them.]

Temperature is remarkably stable, sitting in a band < 1°C wide as the satellite goes in and out of sunlight. The pressure and RH clearly show the effects of the orbital period, as expected; the temperature barely shows it; and the optical readings show no sign of the major temp/illumination swings going on outside having any effect at all on the biology.

My hat’s off to the whole team for an incredibly successful demonstration! This is success.


GeneSat Mission Status Report

Submitted by: Bruce Yost/GeneSat Mission Manager

Mission Status:

We are now in EXPERIMENT phase. The experiment has been activated and data is being downlinked both via the beacon and the primary COM link (2-4 passes/day). We are currently working NO anomalies.

Spacecraft Status:

• All GeneSat systems operating nominally

  • Beacon transmitter functional
  • 2.4 GHz CMD link operational

• Experiment was activated Monday morning at 5:00 AM local.

  • Temperature = 34°C (nominal)
  • RH = 85% (nominal)
  • Pressure = 14.6 – 14.9 psi (nominal)

Ground Segment Status:

• Currently operating from SRI GND station (Stanford foothills)

• Successfully completed 3 contacts this morning

  • o First look experiment data downloaded

• Beacon contest underway

  • o Received >4200 beacon packets from global HAM community to date
  • o Contest will end after Thursday

• Issues

  • o SCU power outage (Tuesday AM)
    • Impact to data distribution, only
    • Work around is to route data via email
    • Overall impact to operations is minimal
    • Cause of outage under investigation – appears to be campus-wide

  • o ARC Firewall reset
    • Occurred during contacts Tuesday AM
    • NO prior notification provided to GeneSat team
    • Prevented transmission of CMDs from secure server to GND station
    • Ops Team implemented work-around using local assets at SRI

Biology/Technology Status:

• First look data indicate all payload systems fully operational

  • Growth curves (OD) indicate growth of bacteria
  • Fluorescent data are within pre-flight expectations
  • Experimental data quality is excellent

SpaceRef staff editor.